OpenStudy (tayfishb):

Find a point c satisfying the conclusion of the MVT for the given function and interval y=x^3 [-4,5]

OpenStudy (tayfishb):

\[f \prime c=\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (tayfishb):

\[\frac{ 5^{3}-(-4^{3}) }{ 5-4 }=189\]

OpenStudy (tayfishb):

\[f \prime (c)=3x ^{2}\]

OpenStudy (tayfishb):

\[3x ^{2}=189\]

OpenStudy (tayfishb):

\[x ^{2}=63\]

OpenStudy (tayfishb):

so logically \[x=\sqrt{63}\]

OpenStudy (tayfishb):

but in the book it says it is equal to \[\pm \sqrt{7}\]

OpenStudy (tayfishb):

???????????

OpenStudy (tayfishb):

@myininaya

OpenStudy (tayfishb):

@brainzonly

Nnesha (nnesha):

\[\frac{ f(5)-f(-4) }{ b-a } = \frac{5^3- (-4)^3} {5-(\color{Red}{-4})}\]

OpenStudy (tayfishb):

oh my goodness thank you soo much im soo dumb sorry

Nnesha (nnesha):

no you're not. these types of mistakes are common 4 calc students ;P